HUBBLE SPACE TELESCOPE DAILY REPORT #929 PERIOD COVERED: 0800 EST Daily Status Report as of 196/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 5055 (A Critical Test of the Galactic Escape Velocity at R(Sun): Cycle 1 Obsv.) This observation was not executed due to the current Safemode event. (See Note 2.4) 1.2 Completed PC 5014 (WF/PC Cycle 2 Calibration: PC Flats Narrow. 2) This observation was not executed due to the current Safemode event. (See Note 2.4) 1.3 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies with the FOC/96) This observation was not executed due to the current Safemode event. (See Note 2.4) 1.4 Completed Two Sets of FOS/RD 4953 (QSO Absorption Systems Snapshot Survey - Part 3 of 3: Cycle 3 High) Both sets of observations were canceled due to the current Safemode event. 1.5 Completed WFC 5013 (WF/PC Cycle 3 Calibration: WFC Narrow. 2) This observation was not executed due to the current Safemode event. (See Note 2.4) 1.6 Safemode Recovery The return to SMS processing and science observations were successfully completed as scheduled at 195/2200Z. 1.7 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.8 Completed FOS/BL 4600 (FUV Spectroscopy fo Io's SO2 Atmosphere and Surface - Cycle 3 High) The Faint Object Spectrograph (Blue Side) was used to make spectral observations of the bright and dark regions on the surface of Jupiter's moon Io. A scan across Io was also made. The objective of this proposal was to further characterize the behavior fo SO2 in Io's atmosphere. The first observation was received and the spectra was displayed with the wavelength scale in real time. A corrective maneuver was successfully uplinked. Post slew analysis indicated that there was a sign ambiguity in the reported wavelength shift, so the target was probably not entirely in the aperture. The pointing correction was only partially successful. In addition the target was not well-centered along the FOS/X-axis. The scan was supposed to cross the disk, parallel to the FOS/BL Y-axis. The tracking slew did nearly parallel the V1 axis, as it should, but the overlight sums were the greatest in the first readouts, instead of starting at small numbers, increasing, and then decreasing. Subsequent observations showed no apparent problems. 1.9 Completed FOS/BL 4952 (QSO Absorption System Snapshot Survey - Part 2 of 2: Cycle 3 High) The Faint Object Spectrograph was used to make an observation of quasar 1257+2840 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. The observations was executed as scheduled, and no problems were reported. 1.10 Completed GHRS/2 4271 (Testing the Star-Disk Connection: C IV and MG II Maps of Accretion Disks) The High Resolution Spectrograph was used to make spectral observations of the dwarf nova system T-Leonis, which usually has a visual magnitude of 15.5 but may brighten to a magnitude of 10.0 for outburst. The OSS monitored the observations and assessed the data quality. A status buffer error message was received that the GHRS carrousel failed to stabilize during observation number six, which occurred during LOS. There was no impact to the science observations. 1.11 Completed PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make a snapshot image of the protoplanetary system HD89353 to study its current morphology. The observation was complete as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 3 Successful : 3 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 12 Successful : 12 2.3 Significant HSTARs: None 2.4 Operations Note: The Health and Safety load was initiated at 194/1600Z, and the return to SMS processing and science observations were successfully completed as scheduled at 195/2200Z. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/RD 4952 (QSO Absorption System Snapshot Survey - Part 2 of 3: Cycle High) 3.2 Complete FOS/BL 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: Cycle High) 3.3 Complete WFC 4599 (Optical Morphology of Radio Quiet, Cluster Galaxies at Redshift Z = 3.4) 3.4 Complete PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.5 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.6 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.7 Complete PC 4809 (Cycle 2: WF/PC PSF Calibration Proposal) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-